OpenStudy (marysammonss):

Paul’s Plumbing charges a flat fee of $90 plus $40/h for a plumbing job. Perfect Pipes charges a flat fee of $60 plus $55/h for a plumbing job. Write an expression to represent the total cost of a plumbing job for Paul’s Plumbing using h to represent the time, in hours, of a plumbing job.

OpenStudy (tgstudios):

x = 90 + 40h would simply be it!

OpenStudy (marysammonss):

@TGstudios can you help me w the rest

OpenStudy (marysammonss):

Write an expression to represent the total cost of a plumbing job for Perfect Pipes using h to represent the time, in hours, of a plumbing job.

OpenStudy (tgstudios):

same, x = 60 + 55h

OpenStudy (tgstudios):

they probably want the other way, x = 55h + 60, I would put that

OpenStudy (tgstudios):

anything else?

OpenStudy (marysammonss):

Find the cost of a 3-hour-long plumbing job at Paul’s Plumbing? Show how you used your expression to find the cost.

OpenStudy (marysammonss):

ur awesome rn lol

OpenStudy (tgstudios):

x = 40(3) + 90

OpenStudy (tgstudios):

would be x = 120 + 90

OpenStudy (marysammonss):

thanksssss thats it (:

OpenStudy (tgstudios):

good luck!

OpenStudy (tgstudios):

fan?